5 elements of fonts

Encoding

OT1, T1, TC1 Family

cmr, ptm, phv, pcr, psy, pzd Series Shape Size

Layout of family(cmr), series(m), shape(n)

0 1 2 3 4 5 6 7 8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚ ‹ ›
"1x “ ” „ « » – —  � ı ȷ ff fi fl ffi ffl
"2x ␣ ! " # $ % & ’  ( ) * + , - .  /
"3x 0 1 2 3 4 5 6 7 8 9 : ; < = > ?
"4x @ A B C D E F G H I J K L M N O
"5x P Q R S T U V W X Y Z [ \ ] ^_
"6x ‘ a b c d e f  g h i j k l m n o
"7x p q r s t u v w x y z { | } ~ -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ đ §
"Ax ă ą ć č ď ě ę ğ ĺ ľ ł ń ň ŋ ő ŕ
"Bx ř ś š ş ť ţ ű ů ÿ ź ž ż ĳ ¡ ¿ £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì Í Î Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ ç è é ê ë ì í î ï
"Fx ð ñ ò ó ô õ ö œ ø ù ú û ü ý þ ß

Layout of family(cmr), series(m), shape(sl)

0 1 2 3 4 5 6 7 8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚ ‹ ›
"1x “ ” „ « » – —  � ı ȷ ff fi fl ffi ffl
"2x ␣ ! " # $ % & ’  ( ) * + , - .  /
"3x 0 1 2 3 4 5 6 7 8 9 : ; < = > ?
"4x @ A B C D E F G H I J K L M N O
"5x P Q R S T U V W X Y Z [ \ ] ^_
"6x ‘ a b c d e f  g h i j k l m n o
"7x p q r s t u v w x y z { | } ~ -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ đ §
"Ax ă ą ć č ď ě ę ğ ĺ ľ ł ń ň ŋ ő ŕ
"Bx ř ś š ş ť ţ ű ů ÿ ź ž ż ĳ ¡ ¿ £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì Í Î Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ ç è é ê ë ì í î ï
"Fx ð ñ ò ó ô õ ö œ ø ù ú û ü ý þ ß

Layout of family(cmr), series(m), shape(it)

0 1 2 3 4 5 6 7 8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹ ›
"1x “ ” „ « » – —  � ı ȷ ff fi fl ffi ffl
"2x ␣ ! " # $ % & ’  ( ) * + ,  - .  /
"3x 0 1 2 3 4 5 6 7 8 9 : ; < = > ?
"4x @ A B C D E F G H I J K L M N O
"5x P Q R S T U V W X Y Z [ \ ] ^_
"6x ‘ a b c d e f  g h i j k l m n o
"7x p q r s t u v w x y z { | } ~ -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ đ §
"Ax ă ą ć č ď ě ę  ğ  ĺ ľ ł ń ň ŋ ő ŕ
"Bx ř ś š ş ť ţ ű ů ÿ ź ž ż ĳ ¡  ¿ £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì Í Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ ç è é ê ë ì  í  î  ï
"Fx ð ñ ò ó ô õ ö œ ø ù ú û ü ý þ ß

Layout of family(cmr), series(m), shape(sc)

0 1 2 3 4 5  6  7  8 9 A B C D E  F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹  ›
"1x “  ”  „ « » – —  � ı  ȷ ff fi fl ffi ffl
"2x ␣ !  " # $ % & ’  ( ) * + ,  -  .  /
"3x 0 1 2 3 4 5 6  7  8 9 :  ;  < = > ?
"4x @ A B C D E F G H I J K L M N  O
"5x P Q R S T U V W X Y Z [  \  ]  ^_
"6x ‘  a b c d e f  g h i  j k l m n  o
"7x p q r s t u v w x y z {  |  }  ~  -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő  Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ  đ  §
"Ax ă ą ć č ď ě ę  ğ ĺ ľ ł ń ň ŋ  ő  ŕ
"Bx ř ś š ş ť ţ ű  ů ÿ ź ž ż  ĳ ¡  ¿  £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì  Í  Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ  SS
"Ex à á â ã ä å æ ç  è é ê ë  ì  í  î  ï
"Fx ð ñ ò ó ô õ ö œ ø ù ú û ü ý  þ  ß

Layout of family(cmr), series(m), shape(ui)

0 1 2 3 4 5  6  7  8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹  ›
"1x “ ” „ « » – —  � ı ȷ ff fi fl ffi ffl
"2x ␣ !  " #  $ %  &  ’  (  ) * +  ,  -  .  /
"3x 0 1 2 3 4 5  6  7  8 9 :  ;  <  =  >  ?
"4x @ A B C D E F  G H I  J K L M N O
"5x P Q R S T U V W X Y Z [  \  ]  ^_
"6x ‘  a b c  d e  f  g  h i  j  k l  m n  o
"7x p q r  s  t u  v  w x y z {  |  }  ~  -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ  Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ  İ  đ  §
"Ax ă ą ć č  ď ě  ę  ğ  ĺ  ľ  ł ń ň ŋ  ő  ŕ
"Bx ř  ś  š  ş  ť  ţ  ű  ů  ÿ ź ž ż ĳ  ¡  ¿  £
"Cx À Á Â Ã Ä Å Æ  Ç È É Ê Ë Ì  Í  Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ  Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ  ç  è é ê ë  ì  í  î  ï
"Fx ð ñ ò ó ô õ  ö  œ  ø ù ú û ü ý  þ  ß

Layout of family(cmr), series(b), shape(n)

0 1 2 3 4 5 6 7 8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚ ‹ ›
"1x “ ” „ « » – —  � ı ȷ ff fi fl ffi ffl
"2x ␣ ! " # $ % &  ’  ( ) * + ,  - .  /
"3x 0 1 2 3 4 5 6 7 8 9 : ; < = > ?
"4x @ A B C D E F G H I J K L M N O
"5x P Q R S T U V W X Y Z [ \ ] ^_
"6x ‘ a b c d e f  g h i j k l m n o
"7x p q r s t u v w x y z { | } ~ -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ đ §
"Ax ă ą ć č ď ě ę ğ ĺ ľ ł ń ň ŋ ő ŕ
"Bx ř ś š ş ť ţ ű ů ÿ ź ž ż ĳ ¡ ¿ £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì Í Î Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ ç è é ê ë ì  í  î ï
"Fx ð ñ ò ó ô õ ö œ ø ù ú û ü ý þ ß

Layout of family(cmr), series(bx), shape(n)

0 1 2 3  4 5  6  7  8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹  ›
"1x “  ”  „  « » – —  �  ı  ȷ  ff fi fl ffi ffl
"2x ␣ !  " # $ % & ’  (  ) * + ,  -  .  /
"3x 0 1 2 3 4 5  6  7  8 9 :  ;  < = > ?
"4x @ A B C D E F G H I J K L M N O
"5x P Q R S T U V W X Y Z [  \  ]  ^_
"6x ‘  a b c d e  f  g  h i  j  k l  m n o
"7x p q r s  t u v  w x y z { |  }  ~  -
"8x Ă Ą Ć Č Ď Ě Ę Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ  đ §
"Ax ă ą ć č ď ě  ę  ğ  ĺ  ľ ł  ń ň ŋ  ő  ŕ
"Bx ř  ś  š  ş  ť  ţ  ű  ů  ÿ ź ž ż ĳ  ¡  ¿  £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì  Í  Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã ä å æ ç  è é ê ë  ì  í  î  ï
"Fx ð ñ ò ó ô õ  ö  œ ø ù ú û ü ý  þ ß

Layout of family(cmr), series(bx), shape(sl)

0 1 2 3  4 5  6  7  8 9 A B C D E F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹  ›
"1x “ ” „ «  » – —  �  ı ȷ ff fi  fl ffi ffl
"2x ␣ !  " #  $ %  &  ’  (  )  * +  ,  -  .  /
"3x 0 1 2 3  4 5  6  7  8 9 :  ;  <  =  >  ?
"4x @ A B C D E F  G H I  J K L M N O
"5x P Q R S T U V W X Y Z [  \  ]  ^_
"6x ‘  a b c  d e  f  g  h i  j  k  l  m n  o
"7x p q r  s  t  u  v  w  x y z  {  |  }  ~  -
"8x Ă Ą Ć Č Ď Ě Ę  Ğ Ĺ Ľ Ł Ń Ň Ŋ  Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ  İ  đ  §
"Ax ă ą ć  č ď ě  ę  ğ  ĺ  ľ  ł ń ň  ŋ  ő  ŕ
"Bx ř  ś  š  ş  ť  ţ  ű  ů  ÿ ź  ž  ż  ĳ  ¡  ¿  £
"Cx À Á Â Ã Ä Å Æ  Ç È É Ê Ë Ì  Í  Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ  Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã  ä å  æ  ç  è  é  ê  ë  ì  í  î  ï
"Fx ð ñ ò ó  ô õ  ö  œ  ø ù ú û ü  ý  þ  ß

Layout of family(cmr), series(bx), shape(it)

0  1  2  3  4  5  6  7  8  9  A B C D  E  F
"0x `  ´  ˆ  ˜  ¨  ˝  ˚  ˇ  ˘  ¯  ˙  ¸  ˛  ‚  ‹  ›
"1x “  ”  „  « » – —  �  ı  ȷ  ff fi  fl ffi ffl
"2x ␣ !  " # $ % &  ’  (  )  * + ,  -  .  /
"3x 0 1 2 3  4 5  6  7  8 9 :  ;  < = > ?
"4x @ A B C D E F  G H I  J K L M N O
"5x P Q R S T U V W X Y Z [  \  ]  ^_
"6x ‘  a  b  c  d  e  f  g  h  i  j  k  l  m n  o
"7x p  q  r  s  t  u  v  w  x  y  z  {  |  }  ~  -
"8x Ă Ą Ć Č Ď Ě Ę  Ğ Ĺ Ľ Ł Ń Ň Ŋ Ő Ŕ
"9x Ř Ś Š Ş Ť Ţ Ű Ů Ÿ Ź Ž Ż Ĳ İ  đ  §
"Ax ă ą  ć  č  ď  ě  ę  ğ  ĺ  ľ  ł  ń ň  ŋ  ő  ŕ
"Bx ř  ś  š  ş  ť  ţ  ű  ů  ÿ  ź  ž  ż  ĳ  ¡  ¿  £
"Cx À Á Â Ã Ä Å Æ Ç È É Ê Ë Ì  Í  Î  Ï
"Dx Ð Ñ Ò Ó Ô Õ Ö Œ Ø Ù Ú Û Ü Ý Þ SS
"Ex à á â ã  ä å  æ  ç  è  é  ê  ë  ì  í  î  ï
"Fx ð ñ ò ó  ô õ  ö  œ  ø ù ú û ü  ý  þ  ß